# Improving the statistical efficiency of cross-conformal prediction

### v1

@@ 4.4 Improving cross-conformal prediction
The improvements proposed in the previous subsections are valid for modified cross-conformal prediction; in particular, the new variants are able to produce smaller prediction sets while preserving the same marginal coverage. Specifically, the marginal coverage does not depend on the number of splits $K$ and the number of observations $n$. When the folds have the same size, the techniques can be used to enhance cross-conformal prediction (Vovk,, 2015): in particular, by examining (6), one can observe that it is possible to improve cross-conformal prediction simply by replacing the threshold $\alpha$ with $\alpha+(1-\alpha)(K-1)/(K+n)$ in the prediction sets defined in (10), (12), and (14).
@@
where $\alpha^{\prime}=\alpha+(1-\alpha)(K-1)/(K+n)$. The marginal coverage of the conformal prediction sets $\hat{C}^{\mathrm{e\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$, $\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ and $\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ is at least $1-2\alpha^{\prime}$.
In practice, when $n\gg K$, the prediction sets $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$ and $\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$ are similar. However, for moderate values of $n$, we will see that the sets defined in (10), (12), and (14) are typically narrower than $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$, even though $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$ assures theoretically a lower coverage guarantee. However, improvements are valid as long as the marginal coverage level $1-2\alpha^{\prime}$ is meaningful, which in practical applications is the most common case. It follows that the improvements are not valid, for example, in the extreme case of leave-one-out conformal prediction (the case $K=n$).
@@ 5 Empirical results
where $\hat{\mu}_{\mathcal{D}}$ is the regression function obtained by applying the regression algorithm on $\mathcal{D}$.
@@ 5.1 Simulation study
where the vector of coefficients is drawn as $\beta=\sqrt{10}\,u$ for a uniform random unit vector $u\in\mathbb{R}^{p}$. Ordinary least squares is employed as regression method (if the linear system is underdetermined, then we take the solution that minimizes the $\ell_{2}$-norm). Formally, given the training data $(X_{i},Y_{i}),i\in[n],$ we estimate the regression function $\hat{\mu}(x)=x^{\top}\hat{\beta}$, where $\hat{\beta}=X_{\mathrm{mat}}^{\dagger}Y_{\mathrm{vec}},$ $Y_{\mathrm{vec}}$ is the response vector, $X_{\mathrm{mat}}$ is the matrix of covariates of dimension $n\times p$ and $\dagger$ denotes the Moore-Penrose inverse. The miscoverage rate equals $\alpha=0.1$, the number of replications (for each $p$) is $1000$ and for each replication, we generate a single test point $(X_{n+1},Y_{n+1})$. The number of folds for cross-conformal prediction and its extensions is $K=5$.
Figure 1: Simulation results, showing the size and coverage of the predictive sets for cross-conformal prediction and its variants. In the left plot, peaks are observed at $404,102,286,100$ and $307$ for mod-cross, e-mod-cross, u-mod-cross, eu-mod-cross and cross, respectively. The smaller sets are often obtained using eu-mod-cross that has coverage between $1-2\alpha$ and $1-\alpha$. The randomized method (u-mod-cross) performs similarly to cross-conformal prediction.
@@
Figure 2: Simulation results, showing the size and coverage of the predictive interval obtained using $4$ methods. In particular, split conformal prediction is trained at levels $\alpha$ and $2\alpha$ and it is compared with mod-cross and eu-mod-cross. The modified cross-conformal prediction method always overcovers and tends to produce large prediction sets. Its exchangeable and randomized variant gives good results in terms of size. When $p\in[25,60]$, the average size of the eu-mod-cross method is smaller than that of the split conformal prediction method trained at level $2\alpha$.
@@
Additional results, comparing our methods with other conformal prediction methods, such as jackknife+ and full conformal prediction, are reported in Appendix D.
@@ 5.2 Real data application
Conformal prediction methods are applied to 10 0001000010\,00010 000 data points randomly sampled without replacement; while other $2500$ observations chosen at random from those not part of the training set are used as the test set. The miscoverage rate is set to $\alpha=0.1$ and the procedure is repeated $20$ times to remove the randomness of the split. The methods used are cross-conformal prediction and its variants (with $K=10$) and split conformal prediction. In particular, split conformal prediction is trained both at levels $\alpha$ and $2\alpha$. The averages over $20$ trials are reported as results in Figure 3 and Table 1.
From Figure 3, it is possible to note that cross-conformal prediction and its modified version give very similar results in terms of size and are usually slightly better than split conformal prediction trained at level $\alpha$. The randomized methods u-mod-cross and eu-mod-cross show a significant improvement in terms of size. The improvement is not as evident for the e-mod-cross method, which turns out to be slightly better than the modified method. The smaller sets are obtained using split conformal prediction trained at level $2\alpha$. The level of coverage of cross-conformal prediction (and mod-cross) is around $1-\alpha$ and the two methods tend to overcover (indeed, they guarantee a miscoverage rate smaller than $2\alpha$). The e-mod-cross method exhibits similar performance to cross-conformal prediction in terms of coverage; while the coverage of u-mod-cross and eu-mod-cross is between the levels $1-2\alpha$ and $1-\alpha$. The coverage of split conformal prediction is essentially equal to the target levels $1-\alpha$ or $1-2\alpha$ (see Appendix A for further details on the coverage of the method).
Additional experiments are provided in Appendix E.
@@
Method mod-cross e-mod-cross u-mod-cross eu-mod-cross cross split split($2\alpha$)
LM 0.903 0.899 0.851 0.858 0.902 0.902 0.800
Lasso 0.902 0.900 0.851 0.849 0.901 0.898 0.801
RF 0.903 0.894 0.853 0.847 0.903 0.900 0.799
Table 1: Empirical coverage for the News Popularity Dataset using different regression algorithms and different conformal prediction methods. Mod-cross and cross have empirical coverage around $1-\alpha$ (while guaranteeing a coverage level of $1-2\alpha)$. The coverage for the randomized methods lies between $1-2\alpha$ and $1-\alpha$. The e-mod-cross variant has coverage $\approx 1-\alpha$.
@@ 6 Discussion
We present new variants of cross-conformal prediction that can achieve smaller prediction sets while maintaining valid coverage guarantees. The achievements are based on recent results on the combination of dependent and exchangeable p-values. In particular, starting from a miscoverage rate equal to $\alpha$, the new methods guarantee a marginal coverage of at least $1-2\alpha$. The same coverage is guaranteed by other methods, such as the jackknife+ by Barber et al., (2021) or multi-split conformal prediction (with threshold set to a half) by Solari and Djordjilović, (2022).
Specifically, similar to cross-validation, the proposed approaches require training the models only $K$ times, unlike $n$ times for the jackknife+ or even potentially an infinite number of times for full conformal prediction (see Table 2). The empirical coverage of the proposed methods usually oscillates between levels $1-2\alpha$ and $1-\alpha$, while for cross-conformal the empirical coverage is usually $\approx 1-\alpha$. As reported in the experimental results, the size of the sets is smaller than that obtained by split conformal prediction and cross-conformal prediction. Since the results depend on randomized or asymmetric combinations of p-values, the size is generally, though not consistently, more variable compared to cross-conformal prediction. In particular, while randomization and asymmetric combination improve efficiency of the prediction sets, they add an extra-layer of randomness to the procedure. The results presented in Sections 4.1, 4.2 and 4.3 can also be extended to the case where the number of observations in the folds are different. Cross-conformal prediction can be improved using the same techniques, as shown in Section 4.4.
In general, the new variants of cross-conformal prediction show good properties in terms of coverage and size of the sets, both in simulations and practical applications.
@@
Table 2: Comparison of the properties of different conformal prediction methods. The first two columns regard the marginal coverage. In particular the theoretical guarantees are valid in finite sample. The last column counts the number of times that the algorithm $\mathcal{A}$ is run on a data set containing $n$ training points, for obtaining a prediction set for a new test point. The parameter $n_{\mathrm{grid}}$ represents the number of different possible $y$ values.
@@ Appendix B Marginal coverage of cross-conformal prediction and connection with CV+
In addition, cross-conformal prediction (Vovk,, 2015) is closely related to $K$-fold CV+ introduced in Barber et al., (2021). In particular, both methods can be used to obtain prediction sets with finite sample coverage guarantees. Cross-conformal prediction is covered in Section 3; here, we introduce CV+ and explain its connection to cross-conformal prediction. In this case as well, the data points are divided into $K$ disjoint folds $I_{1},\dots,I_{K}$ of size $m=n/K$, and $\hat{\mu}_{-I_{k}}$ refers to the regression function trained in $[n]\setminus I_{k}$, $k\in[K]$. The $K$-fold CV+ prediction set is defined as
$\begin{split}C^{\mathrm{CV+}}_{n,K,\alpha}(X_{n+1})=\bigg{[}&-\mathrm{quantile}\left(\left(-\left(\hat{\mu}_{-I_{k(i)}}(X_{n+1})-S_{i}^{\mathrm{CV+}}\right)\right)_{i\in[n]};(1-\alpha)(1+1/n)\right),\\ &\mathrm{quantile}\left(\left(\hat{\mu}_{-I_{k(i)}}(X_{n+1})+S_{i}^{\mathrm{CV+}}\right)_{i\in[n]};(1-\alpha)(1+1/n)\right)\bigg{]},\end{split}$ (20)
where $S_{i}^{\mathrm{CV+}}=|Y_{i}-\hat{\mu}_{-I_{k(i)}}(X_{i})|$, $i\in[n]$, are the residual scores (or absolute residuals). An attractive property of the set in (20) is that it is very interpretable, since it is always an interval (rather than, possibly, a union of intervals). In particular, when $n=K$, it corresponds to the jackknife+ interval by Barber et al., (2021).
@@
As a corollary, it follows that the marginal coverage guarantee in (19) also holds for the CV+ method. This implies that the marginal coverage guarantee for the jackknife+, the case $K=n$, is at least $1-2\alpha$. The empirical coverage often exceeds the stated $1-2\alpha-2/\sqrt{n}$, typically aligning closer to the level $1-\alpha$, and sometimes approaching one. In fact, the value $2\alpha$ can be considered as the worst-case scenario for the method. In fact, under some assumptions about the stability of the prediction algorithm, a modified version of the jackknife+ is shown to have a marginal coverage close to $1-\alpha$. A similar problem is studied in Steinberger and Leeb, (2023), where the authors prove a conditional coverage probability statement for $K$-fold cross-validation (a set similar to the one in (20)) valid under some assumptions on the algorithm and the distribution of the data. We refer to Angelopoulos et al., (2024, Ch. 6) for a detailed discussion of the properties of cross-conformal prediction and jackknife+.
@@ Appendix C Cross-conformal prediction with varying fold sizes
In this section we treat the case where the number of observations in each fold can differ. We consider the same setup as at the beginning of Section 3, and we allow different sizes among the subsets. Let $m_{k}$ denote the number of observations in subset $I_{k}$, $k\in[K]$. By definition, the sum $m_{1}+\dots+m_{K}$ equals the number of observations $n$. In this case, the definition of conformal p-values in (5) change slightly, allowing for dependence on $m_{k}$ in the denominator:
@@
$\mathbb{P}(Y_{n+1}\in\hat{C}_{n,K,\alpha}^{\mathrm{cross}}(X_{n+1}))\geq 1-2\alpha-2(1-\alpha)\frac{1-1/K}{n/K+1}.$
@@
Figure 5: Simulation results, showing the size and the coverage of the predictive intervals for jackknife+, split conformal prediction and full conformal prediction. The eu-mod-cross method is added for comparison. The smaller sets are usually observed by eu-mod-cross conformal prediction. Split conformal prediction and jackknife+ have empirical coverage $\approx 1-\alpha$.

### v2

@@ 4.4 Improving cross-conformal prediction
The improvements proposed in the previous subsections are valid for modified cross-conformal prediction; in particular, the new variants are able to produce smaller prediction sets while preserving the same marginal coverage. Specifically, the marginal coverage does not depend on the number of folds $K$ and the number of observations $n$. When the folds have the same size, the techniques can be used to enhance cross-conformal prediction (Vovk,, 2015): in particular, by examining (6), one can observe that it is possible to improve cross-conformal prediction simply by replacing the threshold $\alpha$ with $\alpha+(1-\alpha)(K-1)/(K+n)$ in the prediction sets defined in (10), (12), and (14).
@@
where $\alpha^{\prime}=\alpha+(1-\alpha)(K-1)/(K+n)$. If data are exchangeable, the marginal coverage of the conformal prediction sets $\hat{C}^{\mathrm{e\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$, $\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ and $\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ is at least $1-2\alpha^{\prime}$.
In practice, when $n\gg K$, the prediction sets $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$ and $\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$ are similar. However, for moderate values of $n$, we will see that the sets defined in (10), (12), and (14) are typically narrower than $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$, even though $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$ assures theoretically a lower coverage guarantee. Clearly, the proposed improvements are valid as long as the marginal coverage level $1-2\alpha^{\prime}$ is meaningful, which in practical applications is the most common case. It follows that the improvements are not valid, for example, in the extreme case of leave-one-out conformal prediction (the case $K=n$).
@@ 5 Empirical results
where $\hat{\mu}_{\mathcal{D}}$ is the regression function obtained by applying the regression algorithm $\mathcal{A}$ on $\mathcal{D}$.
The code to reproduce the experiments is available at github.com/matteogaspa/EffCrossCP.
@@ 5.1 Simulation study
where the vector of coefficients is drawn as $\beta=\sqrt{10}\,v$ for a uniform random unit vector $v\in\mathbb{R}^{p}$. Ordinary least squares is employed as regression method (if the linear system is underdetermined, then we take the solution that minimizes the $\ell_{2}$-norm). Formally, given the training data $(X_{i},Y_{i}),i\in[n],$ we estimate the regression function $\hat{\mu}(x)=x^{\top}\hat{\beta}$, where $\hat{\beta}=X_{\mathrm{mat}}^{\dagger}Y_{\mathrm{vec}},$ $Y_{\mathrm{vec}}$ is the response vector, $X_{\mathrm{mat}}$ is the matrix of covariates of dimension $n\times p$ and $\dagger$ denotes the Moore-Penrose inverse. The nominal miscoverage rate equals $\alpha=0.1$, the number of replications (for each $p$) is $1000$ and for each replication, we generate a single test point $(X_{n+1},Y_{n+1})$. The number of folds for cross-conformal prediction and its extensions is $K=5$.
Figure 1: Simulation results, showing the size and coverage of the predictive sets for cross-conformal prediction and its variants. In the left plot, peaks are observed at $404,102,286,100$ and $307$ for mod-cross, e-mod-cross, u-mod-cross, eu-mod-cross and cross, respectively. The parameter $\alpha$ is set to 0.1. The smaller sets are often obtained using eu-mod-cross that has coverage between $1-2\alpha$ and $1-\alpha$. The randomized method (u-mod-cross) performs similarly to cross-conformal prediction.
@@
Additional results, comparing the proposed variants with other conformal prediction methods, such as jackknife+ and full conformal prediction, are reported in Appendix D.
Figure 2: Simulation results, showing the size and coverage of the predictive intervals obtained using $4$ methods. The parameter $\alpha$ is set to 0.1. Split conformal prediction is trained at levels $\alpha$ and $2\alpha$ and it is compared with mod-cross and eu-mod-cross. The modified cross-conformal prediction method always overcovers and tends to produce large prediction sets. Its exchangeable and randomized variant gives good results in terms of size. When $p\in[25,60]$, the average size of the eu-mod-cross method is smaller than that of the split conformal prediction method trained at level $2\alpha$.
@@ 5.2 Real data application
Conformal prediction methods are applied to 10 0001000010\,00010 000 data points randomly sampled without replacement; while other $2500$ observations chosen at random from those not part of the training set are used as the test set. The miscoverage rate is set to $\alpha=0.1$ and the procedure is repeated $100$ times to remove the randomness of the split. The method used are cross-conformal prediction and its variants (with $K=10$) and split conformal prediction. In particular, split conformal prediction is trained both at levels $\alpha$ and $2\alpha$. The averages over $100$ trials are reported as results in Figure 3 and Table 1.
@@
[…]
LM Mean 0.902 0.896 0.851 0.852 0.901 0.900 0.801
Min 0.884 0.876 0.834 0.826 0.884 0.885 0.777
Max 0.917 0.915 0.874 0.881 0.916 0.916 0.831
Range 0.032 0.039 0.040 0.055 0.032 0.030 0.054
Lasso Mean 0.901 0.896 0.851 0.845 0.900 0.900 0.800
Min 0.888 0.872 0.831 0.826 0.887 0.884 0.772
Max 0.916 0.914 0.870 0.876 0.915 0.915 0.822
Range 0.028 0.042 0.039 0.050 0.028 0.030 0.050
RF Mean 0.903 0.892 0.853 0.847 0.903 0.900 0.801
Min 0.885 0.870 0.832 0.815 0.885 0.882 0.778
Max 0.918 0.915 0.875 0.881 0.917 0.918 0.819
Range 0.032 0.045 0.042 0.066 0.032 0.036 0.042
Table 1: Empirical coverage for the News Popularity Dataset using different regression algorithms and different conformal prediction methods. The $\alpha$-level is set to $0.1$. Mod-cross and cross have empirical coverage around $1-\alpha$ (while guaranteeing a coverage level of at least $1-2\alpha)$. The coverage for the randomized methods lies between $1-2\alpha$ and $1-\alpha$. The e-mod-cross variant has coverage $\approx 1-\alpha$. Coverage variability is reported through the minimum and maximum values across the 100 replications, together with their difference.
From Figure 3, it is possible to note that cross-conformal prediction and its modified version give very similar results in terms of size and are usually slightly better than split conformal prediction trained at level $\alpha$. The randomized methods u-mod-cross and eu-mod-cross show a significant improvement in terms of size. The improvement is not as evident for the e-mod-cross method, which turns out to be slightly better than the modified method. The smaller sets are obtained using split conformal prediction trained at level $2\alpha$. The level of coverage of cross-conformal prediction (and mod-cross) is around $1-\alpha$ and the two methods tend to overcover (indeed, they guarantee a miscoverage rate smaller than $2\alpha$). The e-mod-cross method exhibits similar performance to cross-conformal prediction in terms of coverage; while the coverage of u-mod-cross and eu-mod-cross is between the levels $1-2\alpha$ and $1-\alpha$. Clearly, there exists a direct relationship between set size and coverage, with smaller prediction sets attaining coverage values closer to $1-2\alpha$. The coverage of split conformal prediction is essentially equal to the target levels $1-\alpha$ or $1-2\alpha$ (see Appendix A for further details on the coverage properties of the method). As a final remark, the eu-mod-cross method demonstrates a higher degree of variability in coverage, while the remaining methods exhibit comparable and more stable variability levels.
Additional experiments on real-world datasets are reported in Appendix E.
@@ 6 Discussion
We present new variants of cross-conformal prediction that can achieve smaller prediction sets while maintaining valid coverage guarantees. The achievements are based on recent results on the combination of dependent and exchangeable p-values (Gasparin et al.,, 2025). In particular, starting from a target miscoverage rate equal to $\alpha$, the new methods guarantee a marginal coverage of at least $1-2\alpha$. The same coverage is guaranteed by other methods, such as the jackknife+ introduced by Barber et al., (2021) or multi-split conformal prediction (with threshold set to a half) by Solari and Djordjilović, (2022).
Specifically, similar to cross-validation, the proposed approaches require training the models only $K$ times, unlike $n$ times for the jackknife+ or even potentially an infinite number of times for full conformal prediction (see Table 2). The empirical coverage of the proposed methods usually oscillates between levels $1-2\alpha$ and $1-\alpha$, while for standard cross-conformal prediction the empirical coverage is usually around the nominal $1-\alpha$ level. As reported in the experimental results, the size of the sets is smaller than that obtained by split conformal prediction and cross-conformal prediction. Since the results depend on randomized or asymmetric combinations of p-values, the size is generally, though not consistently, more variable compared to cross-conformal prediction. In particular, while randomization and asymmetric combination improve efficiency of the prediction sets, they add an extra-layer of randomness to the procedure. The results presented in Sections 4.1, 4.2 and 4.3 can also be extended to the case where the number of observations in the folds are different. Cross-conformal prediction can be improved using the same techniques, as shown in Section 4.4.
The question of which variant of cross-conformal prediction to use in practice is a subtle one. If it is crucial to obtain a $1-\alpha$ guarantee against worst-case distributions and unstable algorithms (for example, when downstream decisions critically depend on the provided theoretical guarantees), then it makes sense to run our methods at level $\alpha/2$. On the other hand, if conformal prediction is used merely as “weak guidance” for downstream decisions, meaning the user is willing to tolerate violations of the $1-\alpha$ target, then we recommend running the proposed cross-conformal variants at level $\alpha$ (where the worst-case coverage is $1-2\alpha$, but actual coverage lies between $1-2\alpha$ and $1-\alpha$). Finally, if the situation is intermediate, where conformal prediction is neither applied extremely rigorously nor used as loose guidance, and the goal is to achieve $1-\alpha$ coverage for “typical” datasets and algorithms, while accepting both overcoverage and undercoverage for unusual distributions or algorithms, then (modified) cross-conformal prediction may be the most appropriate choice.
In general, the sets presented in Section 4 show good properties in terms of size and coverage, both in simulations and in applications. The methods can be especially useful in settings where full conformal prediction or jackknife+ are computationally prohibitive.
@@
Table 2: Comparison of the properties of different conformal prediction methods. The first two columns regards the marginal coverage. The theoretical guarantees are valid in finite sample. The last column counts the number of times that the algorithm $\mathcal{A}$ is run on a data set containing $n$ training points, for obtaining a prediction set for a new test point. The parameter $n_{\mathrm{grid}}$ represents the number of different possible $y$ values.
@@ Appendix B Marginal coverage of cross-conformal prediction and connection with CV+
In addition, in a regression setting, cross-conformal prediction (Vovk,, 2015) is closely related to $K$-fold CV+ introduced in Barber et al., (2021). In particular, both methods can be used to obtain prediction sets with finite sample coverage guarantees. Cross-conformal prediction is covered in Section 3; here, we introduce CV+ and explain its connection to cross-conformal prediction. In this case as well, the data points are divided into $K$ disjoint folds $I_{1},\dots,I_{K}$ of size $m=n/K$, and $\hat{\mu}_{-I_{k}}$ refers to the regression function trained using data in $[n]\setminus I_{k}$, $k\in[K]$. The $K$-fold CV+ prediction set is defined as
$\begin{split}\hat{C}^{\mathrm{CV+}}_{n,K,\alpha}(X_{n+1})=\bigg{[}&-\mathrm{quantile}\left(\left(-\left(\hat{\mu}_{-I_{k(i)}}(X_{n+1})-S_{i}^{\mathrm{CV+}}\right)\right)_{i\in[n]};(1-\alpha)(1+1/n)\right),\\ &\mathrm{quantile}\left(\left(\hat{\mu}_{-I_{k(i)}}(X_{n+1})+S_{i}^{\mathrm{CV+}}\right)_{i\in[n]};(1-\alpha)(1+1/n)\right)\bigg{]},\end{split}$ (20)
where $S_{i}^{\mathrm{CV+}}=|Y_{i}-\hat{\mu}_{-I_{k(i)}}(X_{i})|$, $i\in[n]$, are the residual scores (or absolute residuals). An attractive property of the set in (20) is that it is interpretable, since it is always an interval (rather than, possibly, a union of intervals). In particular, when $n=K$, it corresponds to the jackknife+ interval by Barber et al., (2021).
@@
As a corollary, it follows that the marginal coverage guarantee in (19) also holds for the CV+ method. This implies that the marginal coverage guarantee for the jackknife+, the case $K=n$, is at least $1-2\alpha$. The empirical coverage often exceeds the stated $1-2\alpha-2/\sqrt{n}$, typically aligning closer to the level $1-\alpha$, and sometimes approaching one. In fact, the value $2\alpha$ can be considered as the worst-case scenario for the method. Under some assumptions about the stability of the prediction algorithm, a modified version of the jackknife+ is shown to have marginal coverage close to $1-\alpha$. A similar problem is studied in Steinberger and Leeb, (2023), where the authors prove a conditional coverage probability statement for $K$-fold cross-validation (a set similar to the one in (20)) valid under some assumptions on the algorithm and the distribution of the data. We refer to Angelopoulos et al., (2024, Ch. 6) for a detailed discussion of the properties of cross-conformal prediction and jackknife+.
@@ Appendix C Cross-conformal prediction with varying fold sizes
In this Section we treat the case where the number of observation in each fold can differ. We consider the same setup as at the beginning of Section 3, and we allow different sizes among the subsets. Let $m_{k}$ denote the number of observations in subset $I_{k}$, $k\in[K]$. By definition, the sum $m_{1}+\dots+m_{K}$ equals the number of observations $n$. In this case, the definition of conformal p-values in (5) change slightly, allowing for dependence on $m_{k}$ in the denominator:
@@
$\mathbb{P}\big{(}Y_{n+1}\in\hat{C}_{n,K,\alpha}^{\mathrm{cross}}(X_{n+1})\big{)}\geq 1-2\alpha-2(1-\alpha)\frac{1-1/K}{n/K+1}.$
@@
Figure 5: Simulation results, showing the size and the coverage of the predictive intervals for jackknife+, split conformal prediction and full conformal prediction. The eu-mod-cross method is added for comparison and the $\alpha$-level is set to $0.1$. The smaller sets are usually observed by eu-mod-cross conformal prediction. Split conformal prediction and jackknife+ have empirical coverage $\approx 1-\alpha$.